When Sodium yellow light (λ = 5890 A0 ) is used in double-slit experiment, the first-order maximum is observed at a distance of 0.05cm from the central maximum. When the light is placed by one of an unknown wavelength, the third-order maximum occurs at 0.065 cm. what is the wavelength of this later light?
Answer #1

For n-th order maximum,

dy/D = nλ

For first order maximum of sodium light,

d * 0.05 / D = 1 * 5890

For third order maximum of unknown light,

d * 0.065 / D = 3λ

So,

3λ / 5890 = 0.065 / 0.05

=> λ = (0.065 / 0.05) * 5890 / 3 = 2552.3 Ao